Session: Nov 2004 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K. Borkowski at Wed Nov 10 06:40:16 2004) Success :
The source went below the telescope horizon at 19:46 UTC. Also, the FS reported once "2004.314.17:57:02.07?ERROR m5 -900 Check formatter serial number even" (though the number is OK). No other known problems.

Medicina (g.maccaferri at Wed Nov 10 06:42:59 2004) Success :

Onsala (Michael Lindqvist at Wed Nov 10 07:58:31 2004) Success :
Last scan on G29 outside trackining limit. Otherwise no known problems.

Effelsberg (A. Kraus at Wed Nov 10 08:26:58 2004) Minor failures : TSYS
High (bad?) T_sys in channels 1 and 3 until UT1245.

Noto (P.Cassaro at Wed Nov 10 09:39:45 2004) Success :

Westerbork (Tony Foley at Thu Nov 11 08:55:41 2004) Success :
Single dish was used. Please change the coordinates in the vexfile accordingly. Changed the RXG file used (c2.rxg) immediately before this experiment. No known problems

Cambridge (Alastair Gunn at Thu Nov 11 22:25:48 2004) Success :
No known problems.

(Other) (Alastair Gunn at Thu Nov 11 22:29:35 2004) Minor failures :
The Darnhall telescope was used. Lost telescope control at 18:07 UT due to vandalism on the communication line to Darnhall. Telescope did not return to control until after the experiment ended. No other known problems.